# “There Has To Be a Lot That We’re Missing”: Moderating AI-Generated Content on Reddit

### v6

@@ 3. Methods: Interview Study
We chose to conduct semi-structured interviews with Reddit moderators, approaching the research questions qualitatively, to deeply “explore and understand the meaning individuals or groups ascribe to a social or human problem” (creswellResearchDesignQualitative2017a), specifically the newly introduced challenges of AIGC. We identified target communities by looking at their public stances towards AIGC, and recruited participants via outreach and snowball methods.
@@ 3.1. Recruitment
As our research goal was to explore moderator concerns, we sought to speak with moderators of communities who had restricted the use of AIGC. To identify such communities, we turned to public data about subreddits’ community rules. We performed a crawl of the Reddit website in the summer of 2023 and extracted the text-based rules from all subreddits indexed by Reddit’s “Top Communities” list111https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We crawled subreddits and found with non-empty rules and English as a primary language. We then performed a text search on the rules to identify subreddits with rules related to AIGC. While this method may be imprecise, it was sufficient for our purpose, which was to understand broad patterns to guide our recruitment strategy.
@@
In the data that we gathered, 4% of all subreddits had rules about AI, but these rules were much more common in larger subreddits: when we restricted our analysis to top 1% of subreddits according to subscriber count, 20% had rules governing the use of AI. We thus designed a recruitment strategy focused on the largest subreddits, which appear to be the most concerned with the potential impact of AIGC. Many moderators of larger subreddits also moderate small forums and were able to speak to the differences between these types of communities.
@@
We recruited and interviewed participants during the summer of 2023. We manually examined the rules of subreddits in descending order of subscriber count. When we found a subreddit with either an explicit rule governing the use of AI or a rule that might implicitly cover AI use we added it to our sample. We also looked for moderator discussions about AIGC in other public forums, such as the r/ModSupport subreddit, and added those moderators’ subreddits to our sample. We then messaged the moderators of each subreddit in our sample via Reddit’s Modmail222https://support.reddithelp.com/hc/en-us/articles/210896606-What-is-Modmail- feature, sending Modmail recruitment messages to 60 subreddits. We used two additional techniques to recruit participants beyond the largest subreddits. First, we conducted snowball sampling and asked moderators to forward our recruitment message to other potential participants. Second, we posted our recruitment call on several social media platforms. To ensure respondents were actual moderators, we asked for their Reddit username and the subreddit that they moderate, then verified that the users were indeed listed as moderators via their public Reddit user profile. Participants were offered $20 gift-cards as compensation. We conducted interviews on a rolling basis and continued recruiting until we had enough rich and complex data to adequately address our research questions, as per Braun and Clarke’s saturation criteria for Reflexive Thematic Analysis (braunSaturateNotSaturate2021). Overall, we interviewed fifteen participants, meeting or even exceeding the sample size norms of the HCI research community (caine16). These participants collectively moderated over 100 different subreddits with sizes ranging from less than ten members to more than 32 million (see participant information in Table 1).
@@ 3.2. Interviews
We designed an initial interview protocol based on our research questions and updated it as themes emerged during early interviews. Since pseudonymity is an important feature of the Reddit platform, we followed best practices in ethical online communities research (Fiesler24) and took special considerations to protect the privacy of our participants. We did not ask for or record participants’ names, demographic information, or any offline identifiers. As text is the primary mode of interaction on Reddit, we gave participants the opportunity to interview via text-based mediums, which is a common practice in Reddit and Social Computing research (Dimond12; Gilbert20; Jhaver19). Accordingly, participants were asked to choose a medium that was comfortable for them for a 60-minute, synchronous interview: video call, audio call, email exchange, or message exchange via the Reddit platform. Regardless of channel, we posed our research questions one at a time, in real time, and followed up in a semi-structured way. For example, for text-based interviews, we were online at the same time as our participants, exchanging messages in a back-and-forth manner, which allowed us to probe participants’ responses as necessary for deeper insight or clarity. The study was approved by our institution’s IRB. All participants read and accepted a consent agreement permitting their responses to be reproduced in research reports along with the participants’ associated subreddits. To ensure that we reliably represented participants’ views, we contacted participants for feedback and approval on their included quotes before publication, following CSCW community recommendations (mcdonald19). We were able to integrate all participants’ feedback in the final manuscript.
@@
How are moderator and non-moderator community members responding to this new phenomenon?
@@
For interviews conducted via audio and video calls, interview transcripts were generated automatically from recordings. The transcripts were manually reviewed and corrected by the first author. For the four interviews conducted via email and the four conducted via Reddit message, the text was captured in similarly formatted transcript files for analysis. Interviews lasted between 26 minutes and 147 minutes with a mean and median of 71 and 62 minutes, respectively. While we scheduled minutes for the interviews, some oral interviews ended early and some text-based interviews went longer, as participants took their time to respond to our questions: oral interviews had a mean duration of minutes (), compared to a mean duration of minutes () for written interviews. Despite the semi-synchronous nature of the longer written interviews, all yielded meaningful conversations with at least eight back-and-forth question and answer exchanges. Written interviews served as a valuable complement to our oral conversations by providing rich data that was not shared via oral conversation, such as links to example AIGC posts from participants’ communities. Despite the differences between the mediums, both yielded deep and meaningful data that combine to create a more nuanced picture than any single medium could provide.
The first two authors then collaborated to iteratively code the interview content into categories, using an inductive thematic analysis method akin to Braun and Clarke’s Reflexive Thematic Analysis (braun21; Braun19). The two authors began by discussing themes that the first author observed while conducting the interviews and correcting the transcripts. From this conversation, an initial high-level set of codes emerged. Next, the first author open-coded all of the interview transcripts at the sentence-level, applying the initial codes and creating new ones where existing codes were insufficient. The two authors then reviewed the set of codes and refined them by merging similar codes and nesting those with a clear hierarchy. This round of coding did not produce new themes, only new sub-codes, suggesting that the themes were fairly stable. Both authors then independently applied these codes to the same three interview transcripts and compared their results. Neither application had created any new codes and the few cases of disagreement were resolved by further merging similar codes. This process yielded approximately 400 tagged phrases across all transcripts, which were grouped into the three high-level themes that we discuss below.
@@
| 2 | r/food (23M) | 25 | Reddit Message Exchange | Lowered Quality | High | ||||||||
@@
| 5 | r/WritingPrompts (17M) | 3 | Video Call | Fewer Opportunities for Skill Development | High | ||||||||
| 6 | r/politics (8.3M) | 25 | Reddit Message Exchange | Strained Community Relationships | Medium | ||||||||
| 7 | r/CryptoCurrency (6.7M) | 8 | Reddit Message Exchange | Amplified Attacks | Medium | ||||||||
| 8 | r/itookapicture (5.1M) | 2 | Email Exchange | Fewer Opportunities for Skill Development | Medium | ||||||||
@@
| 14 | r/AskHistorians (1.8M) | 1 | Reddit Message Exchange | Less Accurate Information | Medium | ||||||||
@@ 4. Findings
Participants generally used the term “AI-generated” in a pejorative way, but some acknowledged that AI-use falls along a spectrum, from benign to problematic, and shared the subjective factors that they considered when judging AI use. Some participants spoke of the intention of a poster333We use the term poster to refer to both those posting top-level subreddit submissions as well as those posting comments in response to submissions., like a moderator of r/itookapicture444To provide additional context, we introduce quotes with the name of the subreddit that the participant moderates. In the case where the participant moderates several subreddits, we include either the subreddit that the quote is referring to or, for more general quotes, the largest subreddit that the participant moderates. who said: “I don’t mind people using AI to better state their thoughts nor create art as long as it is not being used deceptively.” Others referenced the degree to which a poster relied on AI, like a moderator of r/ChangeMyView who reported being OK with a post made with AI, “if it was sufficiently [a poster’s own] words, just sort of juiced up by ChatGPT.” We heard something similar from a participant who cared that a post was a poster’s “own intellectual contribution or content”:
@@
Other participants took a less nuanced view, like a moderator of r/WritingPrompts, who relayed their community’s official stance: “Let’s be absolutely clear: you are not allowed to use AI in this subreddit, you will be banned”. Despite hearing arguments that, “ChatGPT is another writing tool that authors are using and banning ChatGPT will be the same as banning a spellchecker,” this participant disagreed and categorically rejected all uses of AI, going so far as to say that, “[AIGC] is antithetical to everything that [our] subreddit is about.” The simplicity of this stance has its appeal, as more subjective definitions of AIGC introduce the possibility that community members will disagree about specific examples—an implication which we will revisit in later sections.
@@ 4.1. Findings: Moderator Concerns About AIGC (RQ1)
Our recruitment focused on moderators of communities that are actively restricting the use of AIGC. Accordingly, while our participants offered differing definitions of AIGC, they generally opposed its use in their communities. We group our participants’ reported AIGC concerns into three, often overlapping, categories: content quality concerns, social dynamics concerns, and governance concerns. Despite general opposition, participants also acknowledged some potential benefits of AIGC, detailed below.
@@
Table 2 summarizes the three main categories of concerns that our participants shared. Content quality concerns take issue with the quality of AIGC and worry that it decreases the utility of an online community. Social dynamics concerns focus on how AIGC reduces the social value that an online community provides its members. Governance concerns emphasize the ways in which AIGC makes the job of moderating online communities more difficult. Naturally, these categories overlap and interact, and our participants often described considerations that span multiple categories.
@@ 4.1.1. Content Quality Concerns
First and foremost, moderators were concerned about the quality of AIGC. Nine participants reported that AIGC did not meet the quality standards of their subreddits, like a moderator of r/AskHistorians who said that AIGC, “tries to meet the substance and depth of a typical post…however, there are frequent glaring errors in both style and content.” Participants shared several reasons that they considered AIGC low quality, such as their issues with its style, its tendency to be factually inaccurate, and the belief that it was off-topic.
Several participants reported that there were certain stylistic issues with AIGC that made it low-quality content. We heard from one moderator that AI-generated answers to questions in their community often, “do not address the question being asked…tend to be very general and ‘hedge’ more than a real human.” Other moderators put it more bluntly, like a moderator of r/CryptoCurrency who noted that, “[AIGC] is not enjoyable to read”, or a moderator of r/Eldenring who told us: “[AIGC] is content that provides no value/discussion at all, it’s low effort content basically.” This moderator considered effort to be a signal of value or quality, which we will revisit in the discussion. They went on to share that this sort of low effort content can, “[discourage] users who want to post their own original content”.
Four participants who moderate communities that emphasize the accuracy of information (e.g. r/AskHistorians and r/explainlikeimfive) specifically mentioned their concern about inaccurate AI hallucinations. This concern is captured by a moderator of r/explainlikeimfive who objected to AIGC on the grounds that ChatGPT was a “bullshit generator”, or by another participant who shared that AIGC’s inaccuracy was unacceptable in their communities:
@@
Finally, we heard from several moderators that AIGC was simply off-topic in their communities. For example, a moderator of r/itookapicture shared that the goal of their community was, “sharing great photography and photography techniques,” and accordingly, “preventing the submission of AI generated images and human-created compositions alike are a top priority.”
@@ 4.1.2. Social Dynamics Concerns
Of course, online communities are not only a source of high-quality content, but also of meaningful social interactions. We heard from ten participants who were concerned that AIGC would negatively impact the social dynamics in their communities, reducing the social value that their communities provide to members. Participants mentioned several ways that this could happen, such as by decreasing opportunities for human connection, straining relationships, and violating their communities’ values.
Ten participants reported that AI went against their communities’ purpose, which was to share posts created by humans. These participants mentioned the inherently inhuman nature of AIGC and expressed concern that this would lead to less authentic connection in their communities. These were not necessarily objections to the quality of AIGC, but rather to the process through which it is created:
@@
These participants believed people come to their community to connect with others through dialogue and that AIGC does not meaningfully contribute to—or even detracts from—such conversations. A moderator from r/changemyview summarized this sentiment well: “How can we change your view when you haven’t actually stated your view at all? This is ChatGPT’s view. And if I wanted to, I could go argue with ChatGPT. But that’s not helping here.” We heard something similar from six moderators of communities of practice (e.g r/itookapicture, which is devoted to photography) who felt AIGC did not contribute to the collaborative learning that the community aimed to encourage. These are communities where members seek to develop particular skills by doing, sharing, and receiving feedback from other practitioners. Accordingly, posts produced without using their practice did not provide educational value to the community. This concern was summarized by a moderator of r/WritingPrompts who shared that their community existed to, “[give] writers a chance to practice their craft and practice their skill. And if you are taking away the practice element of that, because all you’re doing is feeding a prompt into ChatGPT…The server is no longer serving its purpose.”
Additionally, we heard concern from participants that AIGC could strain relationships in their community. For example, a moderator of a political subreddit noted that: “the possibility [of AIGC] fuels Redditors accusing other comments of being written by AI (which is a form of incivility).” Beyond incivility, the controversial nature of AIGC could cause discussions to devolve into off-topic arguments:
@@
Finally, some participants took an ideological stance and objected to the use of AI and, by extension, AIGC, because they saw it as incompatible with their communities’ values. For example, we heard from a moderator of a creative subreddit who objected specifically to the way that generative AI models are trained:

### v7

@@ 3. Methods: Interview Study
We chose to conduct semi-structured interviews with Reddit moderators, approaching our research questions qualitatively, to deeply “explore and understand the meaning individuals or groups ascribe to a social or human problem” (Creswell and Creswell, 2017), specifically the newly introduced challenges of AIGC. We identified target communities by looking at their public stances towards AIGC, and recruited participants via outreach and snowball methods.
@@ 3.1. Recruitment
As our research goal was to explore moderator concerns, we sought to speak with moderators of communities who had restricted the use of AIGC. To identify such communities, we turned to public data about subreddits’ community rules. We performed a crawl of the Reddit website in the summer of 2023 and extracted the text-based rules from all subreddits indexed by Reddit’s “Top Communities” list111https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We crawled subreddits and found with non-empty rules and English as a primary language. We then performed a text search on the rules to identify subreddits with rules related to AIGC. Though this method may be imprecise, it was sufficient for our purpose, which was to understand broad patterns to guide our recruitment strategy.
@@
In the data that we gathered, 4% of all subreddits had rules about AI, but these rules were much more common in larger subreddits: when we restricted our analysis to the top 1% of subreddits by subscriber count, 20% had rules governing the use of AI. We thus designed a recruitment strategy focused on the largest subreddits, which appear to be the most concerned about AIGC. Many moderators of larger subreddits also moderate small forums and were able to speak about differences of scale.
@@
We recruited and interviewed participants during the summer of 2023. We manually examined the rules of subreddits in descending order of subscriber count. When we found a subreddit with either an explicit rule governing the use of AI or a rule that might implicitly cover AI use, we added it to our sample. We also looked for moderator discussions about AIGC in other public forums, such as the r/ModSupport subreddit, and added those moderators’ subreddits to our sample. We then messaged the moderators of each subreddit in our sample via Reddit’s Modmail222https://support.reddithelp.com/hc/en-us/articles/210896606-What-s-mod-mail-and-how-do-I-send-a-message feature, sending Modmail recruitment messages to 60 subreddits. We used two additional techniques to recruit participants beyond the largest subreddits. First, we used snowball sampling and asked moderators to forward our recruitment message to other potential participants. Second, we posted our recruitment call on several social media platforms. To ensure respondents were actual moderators, we asked for their Reddit username and the subreddit that they moderate, then verified that they were indeed listed as moderators on their Reddit user profile. Participants were offered $20 gift cards as compensation. We conducted interviews on a rolling basis and continued recruiting until we had enough rich and complex data to adequately address our research questions, as per Braun and Clarke’s saturation criteria for Reflexive Thematic Analysis (Braun and Clarke, 2021b). Overall, we interviewed fifteen participants, meeting or even exceeding the sample size norms of the HCI research community (Caine, 2016). These participants collectively moderated over 100 different subreddits with sizes ranging from less than ten members to more than 32 million (see participant information in Table 1).
@@ 3.2. Interviews
We designed an initial interview protocol based on our research questions and updated it as themes emerged during early interviews. Since pseudonymity is an important feature of the Reddit platform, we followed best practices in ethical online communities research (Fiesler et al., 2024) and took special considerations to protect the privacy of our participants. We did not ask for or record participants’ names, demographic information, or any offline identifiers. As is common in Reddit and Social Computing research (Gilbert, 2020; Jhaver et al., 2019b), we gave participants the opportunity to interview via text-based mediums in order to increase participation and representation. Text is the primary mode of interaction on Reddit, so despite the fact that such interviews may produce shorter exchanges and less rich insights than oral interviews (Bruckman, 2006; Dimond et al., 2012), we believed they would be a valuable supplement to our oral interviews. Accordingly, we allowed participants to choose to interview via synchronous video or audio call, semi-synchronous message exchange on the Reddit platform, or asynchronous email exchange. Regardless of channel, we posed our research questions one at a time and followed up in a semi-structured way. For example, for text-based interviews, we were online at the same time as our participants, exchanging messages in a back-and-forth manner, which allowed us to probe participants’ responses as necessary for deeper insight or clarity.
Our study was approved by the Cornell University IRB. All participants read and accepted a consent agreement permitting their responses and subreddit affiliations to be published in research reports. To ensure that we reliably represented participants’ views, we contacted participants for feedback and approval on their included quotes before publication, following CSCW community recommendations (McDonald et al., 2019). We integrated all participants’ feedback in the final manuscript.
@@
How are moderator and non-moderator community members responding to AIGC?
@@
Transcripts for interviews conducted via audio and video calls were generated automatically from recordings, then manually reviewed and corrected by the first author. For interviews done via email and Reddit message, the text was captured in similarly formatted transcript files for analysis. Interviews lasted between 26 minutes and 147 minutes with a mean and median of 71 and 62 minutes, respectively. While we scheduled minutes for interviews, some oral interviews ended early and some text-based interviews went longer, as participants took their time to respond to our questions: oral interviews had a mean duration of minutes (), compared to a mean duration of minutes () for written interviews. Despite the non-synchronous nature of the text-based interviews, all of these interviews yielded meaningful conversations with at least eight back-and-forth question and answer exchanges. Text-based interviews provided rich data that was not shared in oral interviews, such as links to example AIGC posts from participants’ communities. Each interview medium yielded insights that, when combined, yield a more nuanced picture than any single medium could provide.
The first two authors then collaborated to iteratively code the interview content into categories, using an inductive thematic analysis method inspired by Braun and Clarke’s Reflexive Thematic Analysis (Braun and Clarke, 2021a, 2019). The two authors began by discussing themes that the first author observed during interviews and transcript correction. From this conversation, an initial high-level set of codes emerged. Next, the first author open-coded all of the interview transcripts at the sentence-level, applying the initial codes and creating new ones where existing codes were insufficient. The two authors then reviewed the set of codes and refined them by merging similar codes and nesting those within a hierarchy. This round of coding did not produce new themes, only new sub-codes, suggesting that the themes were fairly stable. Both authors then independently applied these codes to the same three interview transcripts and compared their results. Neither created any new codes and the few cases of disagreement were resolved by further merging similar codes. This process yielded approximately 400 tagged phrases across all transcripts, which were grouped into the three high-level themes that we discuss below.
@@
| 2 | r/food (23M) | 25 | Reddit Messages | Lowered Quality | High | ||||||||
@@
| 5 | r/WritingPrompts (17M) | 3 | Video Call | Less Skill Development | High | ||||||||
| 6 | r/politics (8.3M) | 25 | Reddit Messages | Strained Relationships | Medium | ||||||||
| 7 | r/CryptoCurrency (6.7M) | 8 | Reddit Messages | Amplified Attacks | Medium | ||||||||
| 8 | r/itookapicture (5.1M) | 2 | Email Exchange | Less Skill Development | Medium | ||||||||
@@
| 14 | r/AskHistorians (1.8M) | 1 | Reddit Messages | Less Accurate Information | Medium | ||||||||
@@ 4. Findings
Participants generally used the term “AI-generated” in a pejorative way, but some acknowledged that AI-use falls along a spectrum, from benign to problematic, and shared the subjective factors that they considered when judging AI use. Some participants spoke of the intention of a poster333We use the term poster to refer to both those posting top-level subreddit submissions as well as those posting comments in response to submissions., like a moderator of r/itookapicture444To provide additional context, we introduce quotes with the name of the subreddit that the participant moderates. In the case where the participant moderates several subreddits, we include either the subreddit that the quote is referring to or, for more general quotes, the largest subreddit that the participant moderates. who said: “I don’t mind people using AI to better state their thoughts nor create art as long as it is not being used deceptively.” Others referenced the degree to which a poster relied on AI, like a moderator of r/ChangeMyView who reported being OK with a post made with AI, “if it was sufficiently [a poster’s own] words, just sort of juiced up by ChatGPT.” Another participant expressed similar concern that posts should be a poster’s “own intellectual contribution or content”:
@@
Other participants took a less nuanced view, like a moderator of r/WritingPrompts, who relayed their community’s official stance: “Let’s be absolutely clear: you are not allowed to use AI in this subreddit, you will be banned”. Despite hearing arguments that, “ChatGPT is another writing tool that authors are using and banning ChatGPT will be the same as banning a spellchecker,” this participant disagreed and categorically rejected all uses of AI, going so far as to say that, “[AIGC] is antithetical to everything that [our] subreddit is about.” The simplicity of this stance has its appeal, as more subjective definitions of AIGC introduce the possibility that community members will disagree about specific cases—an implication which we will revisit in later sections.
@@ 4.1. Findings: Moderator Concerns About AIGC (RQ1)
Our recruitment targeted moderators of communities that are actively restricting the use of AIGC. Accordingly, while our participants offered differing definitions of AIGC, they generally opposed its use in their communities. We group our participants’ reported AIGC concerns into three, often overlapping, categories, summarized in Table 2. Content quality concerns (detailed below in Section 4.1.1) take issue with the quality of AIGC and worry that it decreases a community’s utility. Social dynamics concerns (Section 4.1.2) focus on how AIGC reduces the social value that a community provides its members. Governance concerns (Section 4.1.3) emphasize the ways in which AIGC makes the job of moderating online communities more difficult. Naturally, these categories overlap and interact, and participants often described considerations that span multiple. Despite general opposition, participants also acknowledged potential benefits of AIGC (Section 4.1.4).
@@ 4.1.1. Content Quality Concerns
First and foremost, moderators were concerned about the quality of AIGC. Nine participants reported that AIGC did not meet the quality standards of their subreddits, like a moderator of r/AskHistorians who said that AIGC, “tries to meet the substance and depth of a typical post…however, there are frequent glaring errors in both style and content.” Participants shared several reasons that they considered AIGC low quality, such as issues with its style, its tendency to be factually inaccurate, and the belief that it was off-topic.
Several participants reported that certain stylistic issues with AIGC make it low-quality content. One moderator said that AI-generated answers to questions often, “do not address the question being asked…tend to be very general and ‘hedge’ more than a real human.” Other moderators put it more bluntly, like a moderator of r/CryptoCurrency who noted that, “[AIGC] is not enjoyable to read”, or a moderator of r/Eldenring who told us: “[AIGC] is content that provides no value/discussion at all, it’s low effort content basically.” This moderator considered effort to be a signal of value or quality, which we will revisit in the discussion. They went on to share that this sort of low effort content can, “[discourage] users who want to post their own original content”.
Four participants who moderate communities that emphasize the accuracy of information (e.g., r/AskHistorians and r/explainlikeimfive) specifically mentioned their concern about inaccurate AI hallucinations. This concern is captured by a moderator of r/explainlikeimfive who objected to AIGC on the grounds that ChatGPT was a “bullshit generator”, or by another participant who shared that AIGC’s inaccuracy was unacceptable:
@@
Finally, several moderators said that AIGC was simply off-topic in their communities. For example, a moderator of r/itookapicture shared that the goal of their community was, “sharing great photography and photography techniques,” and accordingly, “preventing the submission of AI-generated images and human-created compositions alike are a top priority.”
@@ 4.1.2. Social Dynamics Concerns
Of course, online communities are not only a source of high-quality content but also of meaningful social interactions. Ten participants said that they were concerned that AIGC would negatively impact the social dynamics in their communities, reducing the social value that their communities provide to members. Participants mentioned several ways that this could happen: decreasing opportunities for human connection, straining relationships, and violating their communities’ values.
Ten participants reported that AI went against their communities’ purpose, which was to share posts created by humans. These participants mentioned the inherently inhuman nature of AIGC and expressed concern that this would lead to fewer authentic connections in their communities. These were not necessarily objections to the quality of AIGC, but rather to the process through which it is created:
@@
These participants believed people come to their community to connect with others through dialogue and that AIGC does not meaningfully contribute to—or even detracts from—such conversations. A moderator of r/changemyview summarized this sentiment well: “How can we change your view when you haven’t actually stated your view at all? This is ChatGPT’s view. And if I wanted to, I could go argue with ChatGPT. But that’s not helping here.” Six moderators of communities of practice (e.g r/itookapicture, which is devoted to photography) similarly said that AIGC did not contribute to the collaborative learning that the community aimed to encourage. These are communities where members seek to develop particular skills by doing, sharing, and receiving feedback from other practitioners. Accordingly, posts produced without using their practice did not provide educational value to the community. A moderator of r/WritingPrompts summarized this concern when they shared that their community existed to, “[give] writers a chance to practice their craft and practice their skill. And if you are taking away the practice element of that, because all you’re doing is feeding a prompt into ChatGPT…The server is no longer serving its purpose.”
Additionally, participants shared concern that AIGC could strain relationships in their community. For example, a moderator of a political subreddit noted that: “the possibility [of AIGC] fuels Redditors accusing other comments of being written by AI (which is a form of incivility).” Beyond incivility, the controversial nature of AIGC could cause discussions to devolve into off-topic arguments:
@@
Finally, some participants took an ideological stance and objected to the use of AI and, by extension, AIGC, because they saw it as incompatible with their communities’ values. For example, a moderator of a creative subreddit objected specifically to the way that generative AI models are trained: